An activity on a PERT network has these time​ estimates: optimistic​ = 2, most likely​ = 3, and pessimistic​ = 10. What is its expected activity time and

variance​?

A.

​5; 1.78

B.

​4; 1.78

C.

​5; 10.67

D.

​4; 10.67

E.

None of the above.
Answer #1

Answer: b.4; 1.78

Expected time =(2+4(3)+10)/6

= 2+12+10/6 = 24/6 = 4

Expected time = 4

Standard deviation

=(10-2)/6

=8/6

= 1.3333

Variance =((10-2)/6)2

= 1.3333*1.3333

Variance =1.78